OpenStudy (anonymous):

How to evaluate 3log3(14) without a calculator?

OpenStudy (anonymous):



OpenStudy (amistre64):

what do you mean by "evaluate"?

OpenStudy (anonymous):

There is a rule in log: \[\log(a \cdot b) = \log(a) + \log(b)\]

OpenStudy (anonymous):

You can write 14 as \(7 \times 2\).. \[3 \log_3(7 \times 2) = 3(\log_3(7) + \log_3(2))\] And after that I will have to use calculator, ha ha ha...

OpenStudy (amistre64):

i can think of using the derivative as the slope of a tangent line at some point, and getting approximations

OpenStudy (anonymous):

But I can't..

OpenStudy (amistre64):